# Quasinormal modes and echoes of a double braneworld

### v1

@@ Abstract
In this work, we study the gravitational quasinormal modes and the gravitational echoes of a double braneworld. The double braneworld is a kind of split thick brane, which is crucial for addressing the hierarchy problem in the thick brane scenarios. Using the Bernstein spectral method, direct integration method, and asymptotic iteration method, we calculate the quasinormal mode frequencies of the double brane. We find that the quasinormal spectrum is very different from that of the single brane model, especially the high overtone mode. We also perform numerical evolution to study the time-domain properties of the characteristic modes of the double brane. The results show that when the degree of brane splitting is large, gravitational echoes of oscillation attenuation between sub-branes will appear in the thick brane. Furthermore, different long-lived Kaluza-Klein modes interfere with each other, resulting in a beating effect. Compared to a single brane model, the phenomenon of the split double brane is richer, and the lifetime of the massive Kaluza-Klein graviton of the double brane is longer. These phenomena may have potential phenomenological interest. We hope to detect these extra-dimensional signals in future gravitational wave detectors or accelerators.
@@ I Introduction
Braneworld models have been a topic of extensive research for many years, offering a new perspective on the nature of spacetime and providing a novel approach to address the hierarchy problem between the Planck and Electroweak scales ArkaniHamed:1998rs ; Antoniadis:1998ig ; Randall:1999ee . The Randall-Sundrum (RS) models, which involve branes embedded in a five-dimensional anti-de Sitter spacetime, have been particularly influential in this regard. The RS-I model Randall:1999ee proposes a solution to the hierarchy problem by introducing two branes, while the RS-II model Randall:1999vf extends this concept by pushing one brane to infinity. A remarkable feature of the RS-II model is the recovery of the four-dimensional Newtonian potential, even with an infinite extra dimension. These brane models have found wide-ranging applications in various areas of physics, including black hole physics, particle physics, and cosmology Shiromizu:1999wj ; Tanaka:2002rb ; Gregory:2008rf ; Jaman:2018ucm ; Adhikari:2020xcg ; Geng:2020fxl ; Geng:2021iyq ; Geng:2022dua ; Bhattacharya:2021jrn .
@@ II Braneworld model in general relativity
$A(y)$ $=$ $\delta H(y)=\frac{-\delta}{2}\ln\left(1+\left(\frac{ky}{\delta}\right)^{2s}\right),$ (8)
@@
$V(\varphi)$ $=$ $3k^{2}\left(\frac{2s+4\delta-1}{2\delta}\cos\left(\frac{s\varphi}{\sqrt{3\delta(2s-1)}}\right)^{2}-2\right)$ (10)
×$\times\left(\sin\left(\frac{s\varphi}{\sqrt{3\delta(2s-1)}}\right)\right).$
@@
$\rho(y)$ $=$ $-3e^{-2H(y)}\left(A^{\prime\prime}(y)-A^{\prime}(y)H^{\prime}(y)+2A^{\prime}(y)^{2}\right)$
$=$ $-\frac{3\delta\left(\left(\frac{ky}{\delta}\right)^{2s}+1\right)^{\frac{1}{s}-2}\left(\frac{ky}{\delta}\right)^{2s}\left(2\delta\left(\frac{ky}{\delta}\right)^{2s}-2s+1\right)}{y^{2}}.$
@@
(b)  $s=1$
Figure 4: Plots of the energy density (LABEL:energydenisty).
@@
$Q^{\dagger}Q\tilde{\phi}(z)=\left(-\partial_{z}^{2}+U^{\text{dual}}(z)\right)\tilde{\phi}(z)=m^{2}\tilde{\phi}(z),$ (25)
@@
III.1 QNMs of thick brane
@@
$A(z)$ $=$ $\frac{-1}{2}\ln\left(1+\left(\frac{kz}{\delta}\right)^{2s}\right),$ (27)
@@
$-\partial_{z}^{2}\phi(z)+\left(\frac{3k^{2s}z^{2s-2}\left(2-4s+5k^{2s}z^{2s}\right)}{4\left(k^{2s}z^{2s}+1\right)^{2}}-m^{2}\right)\phi(z)=0.$
@@
$\tilde{\phi}(z)\propto\left\{\begin{aligned} e^{imz},&~{}~{}~{}~{}~{}z\to\infty.&\\ e^{-imz},&~{}~{}~{}~{}~{}z\to-\infty.&\end{aligned}\right.$ (35)
Because the Bernstein spectral method is effective in the compact coordinate domain, but here the extra dimensions are infinite. The Bernstein spectrum method cannot be used directly. We need to transform our coordinates to be finite. So we introduce the transformation $u=\frac{\sqrt{4k^{2}z^{2}+1}-1}{2kz}$. Now, Eq. (III.1) can be rewritten as
$\left(\frac{m^{2}}{k^{2}}-\frac{3\left(\frac{u}{1-u^{2}}\right)^{2(s-1)}\left(\left(\frac{u}{1-u^{2}}\right)^{2s}+4s-2\right)}{4\left(\left(\frac{u}{1-u^{2}}\right)^{2s}+1\right)^{2}}\right)\phi(u)$
@@
$\phi(u)\propto\left\{\begin{aligned} e^{-\frac{im/k}{2u-2}},&~{}~{}~{}u\to 1.&\\ e^{\frac{im/k}{2u+2}},&~{}~{}~{}u\to-1.&\end{aligned}\right.$ (37)
@@
$-\frac{3\left(u^{2}+1\right)^{3}\left(\left(\frac{u}{1-u^{2}}\right)^{2s}+4s-2\right)\left(\frac{u}{1-u^{2}}\right)^{2s}}{u^{2}\left(\left(\frac{u}{1-u^{2}}\right)^{2s}+1\right)^{2}}.$
We use the Mathematica package Fortuna:2020obg which is called SpectralB to obtain the QNMs of the thick brane. The results are compared with those obtained by the direct integration method, which are shown in Tab. 1. The effect of the parameter $s$ on the first three quasinormal frequencies (QNFs) is shown in Fig. 6. From Tab. 1 and Fig. 6, we can see that the real parts of the second and third QNFs and the imaginary parts of the first three QNFs increase with the parameter $s$. The real part of the first QNF first increases with $s$ and then decreases. This shows that the lifetime of the first three QNMs on the brane increases with $s$. It can also be seen that when $s$ changes from 1 to 3, the first three QNFs change more dramatically. Then, with the increase of $s$, the first three QNFs change gently. This is because when $s$ from 1 to 3 the thick brane splits into two sub-brane, and the structure of the thick brane completely changes. Then $s$ increase only gradually changes the thickness of each sub-brane and the distance between sub-branes, so the quasinormal spectrum changes more gently.
$\;\;s\;\;$ $\;\;n\;\;$   Bernstein spectral method         Direct integration method         Time evolution
    $\text{Re}(m/k)$    $\text{Im}(m/k)~{}~{}$ ${}~{}~{}~{}~{}~{}\text{Re}(m/k)$    $\text{Im}(m/k)~{}~{}$ ${}~{}~{}~{}~{}~{}\text{Re}(m/k)$    $\text{Im}(m/k)~{}~{}$
@@
$\;\;s\;\;$ $\;\;n\;\;$   Bernstein spectral method         Asymptotic iteration method
    $\text{Re}(m/k)$    $\text{Im}(m/k)~{}~{}$ ${}~{}~{}~{}~{}~{}\text{Re}(m/k)$    $\text{Im}(m/k)~{}~{}$
@@
$\;\;s\;\;$ $\;\;n\;\;$   Bernstein spectral method         Direct integration method
    $\text{Re}(m/k)$    $\text{Im}(m/k)~{}~{}$ ${}~{}~{}~{}~{}~{}\text{Re}(m/k)$    $\text{Im}(m/k)~{}~{}$
@@ III.2 Time domain: quasinormal and echo waveforms
We have studied the QNMs of the thick brane in the frequency-domain by various methods. These results show how these fluctuations oscillate and decay when the thick brane is perturbed. But this is not complete, only in the frequency-domain research lacks a description of the evolution of these fields waveform, the relative amplitude, etc. Therefore, it is necessary to consider the evolution of the $1+1$ wave equation (18). First, we still consider the case of $\delta=1$, where the effective potential has an analytical expression in $z$ coordinate. We can easily change the coordinates to the light-cone coordinates $u=t+z$ and $v=t-z$ for evolution. Then evolution equation (18) can be written as
@@
We chose the Gaussian wave packet with width $k\sigma=1$ and located at $kv_{c}=20$. We set the parameter $a$ to $a/k=1$. The evolution waveform of Gauss pulse is shown in Fig. 8, which show how the initial wave packet evolves over time at different locations in the extra dimension. It can be seen that after the initial oscillation damping, the evolution of the wave packet tends to be a sinusoid. The frequencies of these final sinusoids does not depend on the extraction point $z_{\text{ext}}$, but the amplitude decreases with $z_{\text{ext}}$. This is due to the existence of a bound zero mode of the thick brane, which does not decay. To illustrate this more clearly, we extract the late maximum amplitude of the evolved waveform at different points and compare it with the analytic zero mode (22), the results are shown in Fig. 9. It can be seen that the profile obtained by the numerical evolution is consistent with the shape of the analytic zero mode.
@@
Figure 9: Compare the zero mode (dot) exited by the Gauss pulse with the analytical zero mode (22) (line).
@@
According to the above results, the split thick brane exhibits a richer phenomenology, including changes in the spectral structure of QNMs, the emergence of long-lived QNMs, echoes, and the beating effect. For the case of $\delta=1$, these phenomena should be extremely weak because the lifetime of the QNMs in this case is still extremely short compared to that of the bound zero mode. For example, in the case of $s=21$ and $\delta=1$, the lifetime of the first QNM is about $10^{-12}$ seconds when $k=10^{-3}$eV. However, for the case of $\delta>1$, as the degree of brane splitting increases, the lifetime of the QNMs of the thick brane will also increase, which may have a considerable degree of impact on the brane. We hope to search for these extra-dimensional messengers in colliders or gravitational wave detectors.
@@ IV Conclusion and discussion
We first review the split thick brane model and the corresponding transverse-traceless tensor fluctuations. We then derive the wave equation and the Schrodinger-like equation along the extra dimension for spin-2 KK gravitons on the brane. Based on the above equations, we study the quasinormal spectrum of the split thick brane in the frequency domain by the Bernstein spectral method, asymptotic iteration method, and direct integration method. The results are shown in Figs. 6, 7, Tabs. 1,  2, and  3. We find that the spectrum for $s>1$ (corresponding to a split brane) is completely different from that for $s=1$ (corresponding to an unsplit brane). However, when $s$ continues to increase, the fundamental mode changes very little, but the high overtone modes change greatly. This indicates that the fundamental mode is relatively stable to the change of potential, while the high overtone modes are extremely sensitive to the change of potential, suggesting that the high overtone modes may be the key to identify the thick brane with little difference in configuration. To complement the results in the frequency domain, we also perform time-domain evolution. The evolution results are shown in Figs. 8 and 10. The frequency of the first QNM obtained by numerical evolution is consistent with the results in the frequency domain. Finally, we study the gravitational echo of the thick brane with $\delta>1$. The results show that echo signals appear for larger values of $\delta$. This is because with the increase of $\delta$, the splitting degree of the thick brane also increases, the effective potential is similar to a resonant cavity, and the gravitational KK mode oscillates and leaks in the quasi-well, forming an echo. In addition, we also found that the different KK modes of the thick brane interfered with each other, resulting in the beating effect. This lays a foundation for further research on the co-evolution of multiple KK modes on the brane. In general, the splitting of the brane enriches the structure of extra dimensions and also leads to more interesting phenomenology. We hope to detect these extra-dimensional signals, especially echo signals, in future gravitational wave detection experiments.

### v2

@@ Abstract
In this work, we study the gravitational quasinormal modes and the gravitational echoes of a double braneworld. The double braneworld is a kind of split thick brane, which is crucial for addressing the hierarchy problem in the thick brane scenarios. Using the Bernstein spectral method, direct integration method, and asymptotic iteration method, we calculate the quasinormal mode frequencies of the double brane. We find that the quasinormal spectrum is very different from that of the single brane model, especially the high overtone mode. We also perform numerical evolution to study the time-domain properties of the characteristic modes of the double brane. The results show that when the degree of brane splitting is large, gravitational echoes of oscillation attenuation between sub-branes will appear in the thick brane. Furthermore, different long-lived Kaluza-Klein modes interfere with each other, resulting in a beating effect. Compared to a single brane model, the phenomenon of the split double brane is richer, and the lifetime of the massive Kaluza-Klein graviton of the double brane is longer. These phenomena may have potential phenomenological interest. We also explore the gravitational wave signature of the KK graviton of the thick brane and find that the violent splitting of the brane may cause the corresponding frequency to fall within the detection range of the high-frequency gravitational wave detector.
@@ I Introduction
Braneworld models have been a topic of extensive research for many years, offering a new perspective on the nature of spacetime and providing a novel approach to address the hierarchy problem between the Planck and Electroweak scales ArkaniHamed:1998rs ; Antoniadis:1998ig ; Randall:1999ee . The Randall-Sundrum (RS) models, which involve branes embedded in a five-dimensional anti-de Sitter spacetime, have been particularly influential in this regard. The RS-I model Randall:1999ee proposes a solution to the hierarchy problem by introducing two branes, while the RS-II model Randall:1999vf extends this concept by pushing one brane to infinity. A remarkable feature of the RS-II model is the recovery of the four-dimensional Newtonian potential, even with an infinite extra dimension. These brane models have found wide-ranging applications in various areas of physics, including black hole physics, particle physics, and cosmology Shiromizu:1999wj ; Tanaka:2002rb ; Gregory:2008rf ; Jaman:2018ucm ; Adhikari:2020xcg ; Geng:2020fxl ; Geng:2021iyq ; Geng:2022dua ; Bhattacharya:2021jrn ; ValeixoBento:2022qca .
@@ II Braneworld model in general relativity
$A(y)$ $=$ $\delta H(y)=\frac{-\delta}{2s}\ln\left[1+\left(\frac{ky}{\delta}\right)^{2s}\right],$ (8)
@@
$V(\varphi)$ $=$ $3k^{2}\left[\frac{2s+4\delta-1}{2\delta}\cos\left(\frac{s\varphi}{\sqrt{3\delta(2s-1)}}\right)^{2}-2\right]$ (10)
$\times\left[\sin\left(\frac{s\varphi}{\sqrt{3\delta(2s-1)}}\right)\right].$
@@
$\rho(y)$ $=$ $-3e^{-2H(y)}\left[A^{\prime\prime}(y)-A^{\prime}(y)H^{\prime}(y)+2A^{\prime}(y)^{2}\right]$ (11)
$=$ $\frac{-3\delta}{y^{2}}\left[\left(\frac{ky}{\delta}\right)^{2s}+1\right]^{\frac{1}{s}-2}\left(\frac{ky}{\delta}\right)^{2s}$
$\times\left[2\delta\left(\frac{ky}{\delta}\right)^{2s}-2s+1\right].$
@@
(b)  $s=3$
Figure 4: Plots of the energy density (11).
@@
$Q^{\dagger}Q\tilde{\phi}(z)=\left[-\partial_{z}^{2}+U^{\text{dual}}(z)\right]\tilde{\phi}(z)=m^{2}\tilde{\phi}(z),$ (25)
@@
III.1 frequcncy domain: QNFs of thick brane
@@
$A(z)$ $=$ $\frac{-1}{2s}\ln\left[1+\left(kz\right)^{2s}\right],$ (27)
@@
$\left[-\partial_{z}^{2}+\frac{3k^{2s}z^{2s-2}\left(2-4s+5k^{2s}z^{2s}\right)}{4\left(k^{2s}z^{2s}+1\right)^{2}}-m^{2}\right]\phi(z)=0.$
@@
[…]
Because the Bernstein spectral method is effective in the compact coordinate domain, but here the extra dimensions are infinite. The Bernstein spectrum method cannot be used directly. We need to transform our coordinates to be finite. Compacting an infinite domain to a finite one by coordinate transformation is a common method in numerical relativity and in solving QNMs of black holes Grandclement:2007sb ; Cho:2011sf ; Jansen:2017oag ; Jaramillo:2020tuu ; Chung:2023wkd . If the form of the compact coordinates is not chosen properly, many problems will arise. Therefore, taking into account the requirements of spectral methods on the form of the equation, we introduce the transformation $u=\frac{\sqrt{4k^{2}z^{2}+1}-1}{2kz}$. This transformation verified to be appropriate in our previous works Tan:2022vfe ; Tan:2023cra ; Tan:2024url . Now, Eq. (III.1) can be rewritten as
$\left\{\frac{m^{2}}{k^{2}}-\frac{3\left(\frac{u}{1-u^{2}}\right)^{2(s-1)}\left[\left(\frac{u}{1-u^{2}}\right)^{2s}+4s-2\right]}{4\left[\left(\frac{u}{1-u^{2}}\right)^{2s}+1\right]^{2}}\right\}\phi(u)$
@@
[…]
@@
$-\frac{3\left(u^{2}+1\right)^{3}\left[\left(\frac{u}{1-u^{2}}\right)^{2s}+4s-2\right]\left(\frac{u}{1-u^{2}}\right)^{2s}}{u^{2}\left[\left(\frac{u}{1-u^{2}}\right)^{2s}+1\right]^{2}}.$
We use the Mathematica package Fortuna:2020obg which is called SpectralB to obtain the QNMs of the thick brane. The results are compared with those obtained by the direct integration method, which are shown in Tab. 1. The effect of the parameter $s$ on the first three quasinormal frequencies (QNFs) is shown in Fig. 6. From Tab. 1 and Fig. 6, we can see that the real parts of the second and third QNFs and the imaginary parts of the first three QNFs increase with the parameter $s$. The real part of the first QNF first increases with $s$ and then decreases. This shows that the lifetime of the first three QNMs on the brane increases with $s$. It can also be seen that when $s$ changes from 1 to 3, the first three QNFs change more dramatically. Then, with the increase of $s$, the first three QNFs change gently. This is because when $s$ from 1 to 3 the thick brane splits into two sub-brane, and the structure of the thick brane completely changes. Then $s$ increase only gradually changes the thickness of each sub-brane and the distance between sub-branes, so the quasinormal spectrum changes more gently. Moreover, to validate our numerical results by the Bernstein spectral method, we performed a backward error analysis on the results. The backward error is defined as Chung:2023wkd
$\Delta m_{n}=|m_{n}(N+i)-m_{n}(N)|.$ (42)
We analyze the backward error of the first QNF for $s=1$ and $s=3$, and the results are shown in Fig. 7. The results show that the truncation error of the first QNF decreases exponentially with the spectral order $N$, consistent with the exponential convergence predicted by the spectral method.
s\;\;s\;\; n\;\;n\;\;   Bernstein spectral method         Direct integration method         Time evolution
    $\text{Re}(m/k)$    $\text{Im}(m/k)$~~ $\text{Re}(m/k)$    $\text{Im}(m/k)$~~ $\text{Re}(m/k)$    $\text{Im}(m/k)$~~
@@
(a)  $s=1$
(b)  $s=3$
Figure 7: The backward error of the first QNF of the thick brane depends on the spectral order NN.
@@
s\;\;s\;\; n\;\;n\;\;   Bernstein spectral method         Asymptotic iteration method
    $\text{Re}(m/k)$    $\text{Im}(m/k)$~~ $\text{Re}(m/k)$    $\text{Im}(m/k)$~~
@@
s\;\;s\;\; n\;\;n\;\;   Bernstein spectral method         Direct integration method
    $\text{Re}(m/k)$    $\text{Im}(m/k)$~~ $\text{Re}(m/k)$    $\text{Im}(m/k)$~~
@@ III.2 Time domain: quasinormal and echo waveforms
We have studied the QNMs of the thick brane in the frequency-domain by various methods. These results show how these fluctuations oscillate and decay when the thick brane is perturbed. But this is not complete, only in the frequency-domain research lacks a description of the evolution of these fields waveform, the relative amplitude, etc. Therefore, it is necessary to consider the evolution of the $1+1$ wave equation (18). First, we still consider the case of $\delta=1$, where the effective potential has an analytical expression in $z$ coordinate. We can easily change the coordinates to the light-cone coordinates $u=t+z$ and v=tzv=t-z for evolution. Then evolution equation (18) can be written as
@@
We chose the Gaussian wave packet with width $k\sigma=1$ and located at $kv_{c}=20$. We set the parameter $a$ to $a/k=1$. The evolution waveform of Gauss pulse is shown in Fig. 9, which show how the initial wave packet evolves over time at different locations in the extra dimension. It can be seen that after the initial oscillation damping, the evolution of the wave packet tends to be a sinusoid. The frequencies of these final sinusoids does not depend on the extraction point $z_{\text{ext}}$, but the amplitude decreases with $z_{\text{ext}}$. This is due to the existence of a bound zero mode of the thick brane, which does not decay. Specifically, according to the analytic solution  (22) of the zero mode and the decomposition  (20), the time-dependent solution of the zero mode is $\phi_{0}(t,z)=\sin(\omega t)\left[1+\left(kz\right)^{2s}\right]^{\frac{-3}{4s}}$. Thus, if we observe at the extra dimensional fixed point, the evolutionary behavior of the zero mode with time is the undecayed sinusoid. To confirm our analysis, we extract the late maximum amplitude of the evolved waveform at different points and compare it with the analytic zero mode (22), the results are shown in Fig. 10. It can be seen that the profile obtained by the numerical evolution is consistent with the shape of the analytic zero mode. This confirms our previous analysis and strengthens the confidence of our numerical evolutionary results.
@@
Figure 10: Compare the zero mode (dot) exited by the Gauss pulse with the analytical zero mode ($2$) (line).
@@
According to the above results, the split thick brane model exhibits a rich phenomenology, including changes in the spectral structure of QNMs, the emergence of long-lived QNMs, echoes, and the beating effect. We also consider how the thick brane affects the scales involved in gravitational wave detection. According to Eq. (16), the equation of motion for $h_{\mu\nu}$ on the brane is given by:
$\square^{(4)}h^{n}_{\mu\nu}-m_{n}^{2}h^{n}_{\mu\nu}=0,$ (46)
where $m_{n}^{2}=\omega^{2}-a^{2}$ is the mass of each KK mode. The zero mode ($n=0$) corresponds to massless gravitons, while higher modes ($n>0$) represent massive spin-2 mode towers.
When an event (such as a black hole merger) produces gravitational waves, it excites $h_{\mu\nu}(x^{\mu};z)$, which decomposes into infinite mode towers on the brane. The zero mode, with $m_{0}=0$, corresponds to gravitational waves with frequencies associated with the source event. The higher modes, with $m_{n}>0$, have frequencies given by $\omega_{n}=m_{n}$, neglecting the effect of $a$. If the KK mode masses are high enough, their frequencies will lie outside the range covered by current gravitational wave detectors, making only the zero mode accessible. However, if the KK mode masses are sufficiently low, their frequencies may fall within the range of current or future experiments. In this case, one might hope to detect a characteristic signature of extra dimensions: a tower of modes with a constant frequency gap $\Delta\omega=m_{1}$. In Tab. 4, we list the frequency ranges and corresponding $m_{1}$ of current or under construction gravitational wave detectors such as Lisa Lisa:2017 , LIGO-Virgo LIGOScientific:2016aoc , and The CERN Axion Solar Telescope (CAST) Ejlli:2019bqj ; Aggarwal:2020olq .
Detector $f_{\text{GW}}$(Hz) $m_{1}$(eV)
LISA  $10^{-4}-10^{0}$ $10^{-31}-10^{-27}$
LIGO-Virgo  $10^{1}-10^{4}$ $10^{-26}-10^{-23}$
CAST  $10^{18}-10^{19}$ $10^{-9}-10^{-8}$
Table 4: The frequency range of the different gravitational wave detectors and the relevant mass range of the gravitational wave corresponding to the massive spin-2 KK mode excitations with mass $\text{Re}(m_{1})$.
In addition to the frequency, the lifetime of the massive KK graviton on the brane, obtained from the imaginary part of the QNM frequency, is crucial for its detectability. Only long-lived KK gravitons can travel far enough to be detected by gravitational wave detectors. Table 5 presents the prospect of KK graviton detection in various brane world scenarios. It can be seen that, for the RS-II thin brane and thick branes with small $s$ and $\delta$, the first QNM frequencies are extremely high, and the lifetimes are very short, making their detection challenging with low- and medium-frequency gravitational wave detectors. However, for thick branes with a large degree of splitting, the first QNM frequency may fall within the detection range of high-frequency gravitational wave detectors such as CAST. Therefore, compared with thin brane or single thick brane, thick branes with sub-brane have richer properties and better observation prospects, we hope to search for these extra-dimensional messengers in high-frequency gravitational wave detectors.
$m_{1}/k$ $\omega_{1}$(Hz) lifetime (s)
RS-II brane   0.4190.577i-0.419-0.577i $10^{22}$   $10^{-13}$
$\delta=s=1$  0.997020.52636i0.99702-0.52636i $10^{23}$   $10^{-13}$
$\delta=3,s=5$  0.450050.01662i0.45005-0.01662i $10^{22}$   $10^{-11}$
$\delta=5,s=35$  0.285640.00180i0.28564-0.00180i $^{22}   1010$$^{-10}$
$\delta=10,s=35$  0.150480.00013i0.15048-0.00013i $10^{22}$   $10^{-9}$
$\delta=20,s=35$  0.077340.000012i0.07734-0.000012i $10^{21}$   $10^{-8}$
Table 5: The frequency and lifetime of the first QNM for branes with different models and parameters. The extra dimension scale $k$ is set to $k=10^{-3}$eV.
@@ IV Conclusion and discussion
We first review the split thick brane model and the corresponding transverse-traceless tensor fluctuations. We then derive the wave equation and the Schrodinger-like equation along the extra dimension for spin-2 KK gravitons on the brane. Based on the above equations, we study the quasinormal spectrum of the split thick brane in the frequency domain by the Bernstein spectral method, asymptotic iteration method, and direct integration method. The results are shown in Figs. 6, 8, Tabs. 1,  2, and  3. We find that the spectrum for $s>1$ (corresponding to a split brane) is completely different from that for $s=1$ (corresponding to an unsplit brane). However, when $s$ continues to increase, the fundamental mode changes very little, but the high overtone modes change greatly. This indicates that the fundamental mode is relatively stable to the change of potential, while the high overtone modes are extremely sensitive to the change of potential, suggesting that the high overtone modes may be the key to identify the thick brane with little difference in configuration. To complement the results in the frequency domain, we also perform time-domain evolution. The evolution results are shown in Figs. 9 and 11. The frequency of the first QNM obtained by numerical evolution is consistent with the results in the frequency domain. Finally, we study the gravitational echo of the thick brane with $\delta>1$. The results show that echo signals appear for larger values of $\delta$. This is because with the increase of $\delta$, the splitting degree of the thick brane also increases, the effective potential is similar to a resonant cavity, and the gravitational KK mode oscillates and leaks in the quasi-well, forming an echo. In addition, we also found that the different KK modes of the thick brane interfered with each other, resulting in the beating effect. This lays a foundation for further research on the co-evolution of multiple KK modes on the brane. In general, the splitting of the brane enriches the structure of extra dimensions and also leads to more interesting phenomenology. We hope to detect these extra-dimensional signals, especially echo signals, in future gravitational wave detection experiments.